# library.bib

### v1

@@ Abstract
Machine learning methods often struggle with real-world applications in science and engineering due to limited or low-quality training data. In this work, the example of groundwater flow with heat transport is considered; this corresponds to an advection-diffusion process under heterogeneous flow conditions, that is, spatially distributed material parameters and heat sources. Classical numerical simulations are costly and challenging due to high spatio-temporal resolution requirements and large domains. While often computationally more efficient, purely data-driven surrogate models face difficulties, particularly in predicting the advection process, which is highly sensitive to input variations and involves long-range spatial interactions. Therefore, in this work, a Local-Global Convolutional Neural Network (LGCNN) approach is introduced. It combines a lightweight numerical surrogate for the transport process (global) with convolutional neural networks for the groundwater velocity and heat diffusion processes (local). With the LGCNN, a city-wide subsurface temperature field is modeled, involving a heterogeneous groundwater flow field and one hundred groundwater heat pump injection points forming interacting heat plumes over long distances. The model is first systematically analyzed based on random subsurface input fields. Then, the model is trained on a handful of cut-outs from a real-world subsurface map of the Munich region in Germany, and it scales to larger cut-outs without retraining. All datasets, our code, and trained models are published for reproducibility.
@@ 1 Introduction
Many real-world systems, across environmental engineering, geoscience, and biomedicine, feature transport phenomena governed by coupled advection–diffusion processes, e.g., pollutant transport through air or water, chemical reactions, or heat transport in flowing media. In such settings, interactions can span large distances, long timescales and many components.
Datasets based on physical measurements are often limited by cost, logistics, or experimental constraints. While numerical simulations can address these limitations to some extent, they become impractically slow or computationally expensive as domain size or resolution increases. Machine learning (ML) models offer fast predictions but typically require large training datasets and often struggle to scale to larger domains. This trade-off underscores the need for data-efficient surrogates that can capture both local and long-range interactions, and scale to arbitrary domain sizes and numbers of interacting components. To this end, we introduce the Local–Global Convolutional Neural Network (LGCNN), a scalable, physics-inspired approach that leverages the structure of advection–diffusion problems by decoupling local and global transport processes.
We apply this approach to the climate change-relevant application of modeling city-wide subsurface temperature fields driven by heat and mass injections of dozens of open-loop groundwater heat pumps (GWHP)[pophillat2020impact, heatpumps, hahnlein2013] in the region of Munich, Germany[geokw22]. A city-wide optimization of the positions of all GWHPs requires calling a prediction model dozens of times with slightly adapted GWHP positions to minimize negative interactions between different heat plumes and pumps. For this scenario, our model should predict the temperature field $T$ in an arbitrarily large domain based on the inputs of the subsurface parameters of a hydraulic pressure gradient $\nabla p$ and a heterogeneous subsurface permeability field $k$, and arbitrary heat pump locations $i$. Labels are generated by expensive simulations, but only for a small number of cases and entirely offline.
As sketched in Figure 1, the LGCNN uses CNNs where they excel, and simple numerical surrogates where CNNs struggle: a first CNN predicts the velocities $\vec{v}$ from local relations of the inputs $p$, $k$ and $i$ (Step 1); a fast numerical component calculates the streamlines $\vec{s}$ based on this (Step 2); a second CNN calculates the temperature distribution $T$ in the whole domain based on all prior inputs and outputs (Step 3). Since Step 2 covers all non-local effects, Step 3 can again take advantage of CNNs to learn local patterns.
@@
We outline related work in Section 2, then define our datasets and metrics in Section 3. In Section 4, we benchmark purely data-driven models and show their performance and limitations on this low-data challenge. To address data scarcity, we introduce LGCNN in Section 5. Section 6 emphasizes generalization and scalability from a single datapoint, while Section 7 demonstrates transfer to real-world subsurface parameters extracted from measurement maps such as in Figure 1.
@@
Our method strongly reduces data requirements for modeling complex scenarios, increases reliability, and ensures scalability to larger domains by a domain-specific modular design of physics-aided machine learning for an input-sensitive scenario with long-distance effects. Subsurface temperature prediction serves as our example application, with the approach expected to generalize to related domains.
@@
The current approach is limited to two-dimensional and steady-state datasets.
2 Related Work
Deep learning surrogates in scientific modeling can be broadly categorized into purely data-driven architectures, physics-informed or domain-structured models, and highly problem-specific hybrid approaches. In this section, we summarize the developments in each category and existing efforts in GWHP modeling.
Convolutional Neural Networks (CNNs), particularly UNet [CNN15] variants, are widely used for spatial prediction tasks due to their locality-aware architecture of moving-window-like kernels. Applications include flow around airfoils [thuerey_airfoils] and various engineering systems [Jhaveri2022A]. For general reviews on ML in scientific and real-world engineering contexts, see [sharma2021machine, sarker2021machine, angra2017machine].
To improve generalization and reduce data requirements, domain-structured models embed physical priors or symmetries into their architecture. This includes Fourier Neural Operators (FNOs) for efficient multi-scale modeling [fno, choi2024applications], thermodynamics-preserving networks [hernandez2021structure], and rotation-equivariant CNNs (ECNNs) [weiler2023ECNNs]. Physics-Informed Neural Networks (PINNs) go further by replacing data loss with physics-based constraints, allowing for training without labeled data [PINN19, cuomo2022scientific], and have been applied to fluid dynamics and inverse problems [pinn_rao, pinn_sun, cai_physics-informed_2021].
Hybrid approaches combine architectures and integrate domain or physics knowledge to improve general models, for instance, in low-data regimes. Examples include staged CNNs for atmospheric plume dispersion [fernandez2024staged], physics-informed FNOs for traffic flow [thodi2024fourier], CNN–MLP hybrids for optimizing positions of oil wells [yousefzadeh2025determination], and physics-guided CNNs for seismic response prediction [zhang2020physics].
Deep Learning for GWHP Modeling
In the context of groundwater modeling with heat transport, recent work focuses on isolated or pairwise interacting GWHPs. Most use UNet-based architectures, optionally with physics-loss terms [davis2023deep, pelzer2024, Stefania]. Although effective in simple, homogeneous aquifers, these models rely on large training datasets or simulated inputs and have not been scaled to city-wide domains with many pumps. UNets are the dominant architecture because of their strong performance on spatial data, making them the baseline for comparison.
We briefly evaluated PINNs and FNOs as alternatives. PINNs, while promising for low-data fluid dynamics tasks [pinn_rao, pinn_sun, cai_physics-informed_2021, pdebench], struggled with complex scenarios (cf. [krishnapriyan2021characterizing]) such as heterogeneous media and discontinuous source terms in preliminary tests. FNOs, though capable of modeling local and global dependencies, struggled with large domain sizes due to high memory requirements. Additionally, they performed poorly in the presence of multiple sources, consistent with [liu2024neural].
@@ 3 Datasets and Metrics
Inputs for the neural networks consist of a heterogeneous permeability field $k$, an initial hydraulic pressure field $p$, and a one-hot-encoded field of heat pump positions $i$. The (interim) labels of velocity $\vec{v}$ and temperature $T$ fields are simulated, in our case with Pflotran [pflotran], until a quasi-steady state is reached after $\approx$27.5 years simulated time [bawue_arbeitshilfe]. All data are normalized to $[0,1]$, and stored in PyTorch format for training.
We generate two types of datasets, one is based on synthetic, the other on real permeability fields $k$. Both cover a $12.8\times 12.8$ km2 domain with 2 560×2 $\times 2\,560$ 2 560 cells. The baseline dataset uses Perlin noise [perlin] to generate random, heterogeneous $k$. Three simulations (3dp) with different fields for $i$ and $k$ are run, generating one datapoint each for training, validation, and testing, plus a simulation 4$\times$ larger to assess scalability. For training vanilla neural networks in Section 4, we generate an additional dataset of 101 datapoints (101dp), split into 73:18:10 for train:val:test. Runtimes are $\approx$27 hours per simulation and 123 hours for the larger domain.
The more realistic dataset builds on $k$ fields cut from maps of borehole measurements in the Munich region[lfu_gepo]. This dataset consists of four simulations (three for training and one for validation) and one larger simulation for testing scalability. Due to constraints of the available measurements, see Figure 1, the large-scale simulation extends only in length, resulting in a rectangular domain twice the length. Runtimes range from 38 to 91 hours (average 58 hours), and 134 hours for the scaling-test domain. Variation across the dataset stems from different heat pump placements in $i$ and from the specific regions where $k$ is extract from the available measurement domain.
For detailed information on simulation procedures and hydro-geological parameters, see Section A.1; for hardware specifications see Section A.4.
@@
We evaluate model accuracy per output dimension ($v_{x},v_{y},T$) separately using Mean Absolute Error (MAE) [naser2023error], Mean Squared Error (MSE) [naser2023error], Maximum Absolute Error ($L_{\infty}$) Structural Similarity Index Measure (SSIM) [ssim], and application-driven metrics of Percentage Above Threshold (PAT) and Visual Assessment. PAT measures the percentage of cells where the absolute error of the predicted temperature exceeds the threshold of $0.1$C, corresponding to measurement precision [kabellichtlot], and is only applicable to temperature predictions. All metrics other than SSIM are applied after re-normalization to the original data ranges to obtain physically meaningful results in [°C] or [meters/year].
4 Application of Purely Data-Driven Approaches
In this section, we present the performance of two purely data-driven approaches for predicting $T$ from inputs of $p,k,i$: a UNet [CNN15] and the domain-decomposition-based UNet variant, called DDUNet [DDUNet].
Methods
The fully convolutional vanilla UNet has an encoder-decoder architecture with skip connections between layers of equal spatial dimensions. Convolutional networks excel in extracting local dependencies from spatial data with locally repeating patterns. Hence, in theory, one datapoint covering a large domain could be enough to generalize, if only local dependencies occur. Global patterns would require a large receptive field and, hence, more data. The DDUNet is designed to handle large-scale domains under GPU memory constraints. It combines a UNet architecture with domain decomposition principles by partitioning the spatial domain into subdomains that are distributed across multiple GPUs. We use a decomposition into $2\times 2$ subdomains. [^1] Conceptually, DDUNet reduces to a standard UNet when trained with a single subdomain. Global context is communicated between subdomains via coarse feature maps through a learned communication network, maintaining spatial consistency and providing a large receptive field while reducing memory demands. The hyperparameters of the models are optimized according to Section A.2.1.
Experiments
Experiments for the vanilla UNet are conducted on the baseline datasets with 3dp and 101dp, those with DDUNet directly on the 101dp dataset, see Section A.3. Both models are applied to the test of 3dp and to the scaling datapoint, for direct comparability to LGCNN. UNet exhibits poor performance when trained on the limited 3dp dataset (UNet3dp), as demonstrated in both Figure 4 and Table 1. Similar performance would be expected for the DDUNet when applied to this sparse dataset. When trained on 101dp, however, both demonstrate robust predictive and scaling capabilities. The advantage of DDUNet over vanilla UNet is a reduction of memory usage and runtime, see section A.3.

### v2

# Resolving Extreme Data Scarcity by Explicit Physics Integration: An Application to Groundwater Heat Transport

@@ Abstract
Real-world flow applications in complex scientific and engineering domains, such as geosciences, challenge classical simulation methods due to large spatial domains, high spatio-temporal resolution requirements, and potentially strong material heterogeneities that lead to ill-conditioning and long runtimes. While machine learning–based surrogate models can reduce computational cost, they typically rely on large training datasets that are often unavailable in practice. To address data-scarce settings, we revisit the structure of advection–diffusion problems and decompose them into multiscale processes of locally and globally dominated components, separating spatially localized interactions and long-range effects. We propose a Local–Global Convolutional Neural Network (LGCNN) that combines a lightweight numerical model for global transport with two convolutional neural networks addressing processes of a more local nature. We demonstrate the performance of our method on city-scale geothermal heat pump interaction modeling and show that, even when trained on fewer than five simulations, LGCNN generalizes to arbitrarily larger domains, and can be successfully transferred to real subsurface parameter maps from the Munich region, Germany.
Machine Learning, ICML
@@ 1 Introduction
Many learning problems in scientific and engineering domains (Sharma et al., 2021; Sarker, 2021; Angra & Ahuja, 2017; Jhaveri et al., 2022) are governed by underlying physical processes. Yet they challenge classical numerical simulations due to large spatial domains, high resolution requirements, and potential ill-conditioning. In addition, many applications require repeated evaluations, for example for uncertainty quantification, design optimization, and real-time control. From a machine learning (ML) perspective, such problems are attractive targets for surrogate modeling, as trained models can dramatically reduce inference cost, even if offline training cost is substantial.
However, in real-world advection-diffusion problems, standard ML approaches face two central difficulties (Dietrich & Schilders, 2025). First, training data are often scarce: measurements are indirect or sparse, and generating labeled data via high-fidelity simulations is computationally expensive. Second, the underlying dynamics exhibit multiscale behavior with both short-range local interactions and long-range global dependencies, making it difficult for data-driven models to generalize in data-scarce regimes. Convolutional architectures such as UNets (Ronneberger et al., 2015a) effectively model local spatial patterns (Thuerey et al., 2020; Jhaveri et al., 2022), but representing non-local interactions requires deep receptive fields and large training datasets. Operator learning approaches, such as Fourier Neural Operators (FNOs) (Li et al., 2020), address long-range dependencies more explicitly in the frequency domain (Choi et al., 2024), but inherently introduce additional assumptions, including periodic boundary conditions, homogeneous coefficients, and fixed domain sizes, which limit their applicability in many real-world settings.
Physics-informed ML aims to mitigate data scarcity by embedding physical constraints or priors into the learning process. Physics-Informed NNs (PINNs) (Dissanayake & Phan-Thien, 1994; Lagaris et al., 1998; Raissi et al., 2019), for instance, replace or augment data losses with squared differential equation residuals, enabling learning with limited labeled data. Despite notable progress (Cuomo et al., 2022; Rao et al., 2020; Sun et al., 2020; Grimm et al., 2025; Wen et al., 2022; Cai et al., 2021), such methods remain at an early stage of methodological maturity and often struggle in complex, high-dimensional scenarios, for example by converging to trivial solutions (Krishnapriyan et al., 2021).
In our approach, rather than increasing architectural complexity, we build on a simple, well-suited base architecture and exploit the physical structure, i.e., of the underlying advection–diffusion system, explicitly. Such systems naturally decompose into processes governed by predominantly local interactions, such as diffusion acting at neighborhood scales and inferable from local information, as well as global transport effects, such as long-range advection, that require propagating information across large parts of the domain. Building on this observation, we propose the Local–Global CNN (LGCNN), a modular architecture that separates local learning from explicit global transport modeling. LGCNN combines CNNs that are well-suited for locally dominant, diffusion-driven processes with a lightweight numerical surrogate that captures long-range transport effects. By inserting this inexpensive, physically motivated component at a critical point in the inference pipeline, LGCNN captures long-range dependencies without relying on deep architectures, large receptive fields, or extensive training data. In contrast to physics-informed approaches that softly impose physical constraints through loss terms, such as PINNs, LGCNN handles global transport at inference time via a dedicated, non-learnable module, leading to improved data efficiency and robustness.
We evaluate LGCNN on the task of predicting city-scale subsurface temperature fields induced by one hundred interacting groundwater heat pumps in the Munich region, Germany (Zosseder et al., 2022). This problem provides a challenging testbed for data-efficient learning, as the temperature field depends on heterogeneous subsurface parameters, arbitrary source configurations, and long-range advective transport over several kilometers and decades. We show that, even when trained on fewer than five simulated samples, LGCNN generalizes to much larger spatial domains without retraining. Our approach also transfers from synthetic inputs to real subsurface parameter maps derived from borehole measurements.
@@
Formally, the learning task is to predict a temperature field $T$ over arbitrarily large domains from heterogeneous permeability fields $k$, hydraulic pressure gradients $\nabla p$, and heat pump locations $i$. Training labels are generated offline using expensive numerical simulations for only a small number of configurations. As illustrated in Figure˜1, LGCNN addresses this setting via a three-stage pipeline: (1) a CNN predicts the velocity field $\vec{v}$ from local relations between $p$, $k$, and $i$; (2) a cheap numerical surrogate computes streamlines $\vec{s}$ to model global transport; and (3) a second CNN predicts the temperature field $T$ conditioned on inputs and intermediate representations. This explicit separation of local learning and global transport is key to the scalability of the approach.
In groundwater heat pump (GWHP) modeling, existing work (Davis et al., 2023; Pelzer & Schulte, 2024; Scheurer, 2021) focuses on isolated or pairwise pump interactions, often using UNet-based architectures with or without physics-based regularization. These methods typically require substantial training data and have not been demonstrated to scale to city-wide domains with many interacting sources and heterogeneous aquifers.
@@
We introduce a modular, physics-inspired architecture that captures complex, multiscale dynamics by combining local convolutional feature learning with explicit long-range transport modeling. To avoid deep architectures and large datasets for modeling nonlocal interactions, LGCNN enforces global coupling through inexpensive numerical transport modules. This structural inductive bias enables inherent generalization and scalable inference across domains and problem sizes from very limited data, i.e., less than five simulations. Compared to classical high-fidelity simulations, our approach achieves an inference-time speedup of approximately 2,000 times.
@@
The current implementation targets a two-dimensional steady-state setting and assumes weak or spatially localized temperature-induced flow perturbations. While this work is restricted to this regime, the proposed framework is easily extendable to three-dimensional and transient problems. Empirical validation of these extensions is left to future work due to the lack of suitable training data.
@@ 2 Datasets and Metrics
Inputs for the NNs consist of a heterogeneous permeability field $k$, an initial hydraulic pressure field $p$, and a one-hot-encoded field of heat pump positions $i$ of a constant injection rate. The (interim) labels of velocity $\vec{v}$ and temperature $T$ fields are simulated with Pflotran (Lichtner et al., 2015a) until a quasi-steady state is reached after $\approx$27.5 years simulated time (Umweltministerium Baden-Württemberg, 2009).
[…]
The more realistic dataset builds on $k$ fields cut from maps of borehole measurements in the Munich region, Germany (Bayerisches Landesamt für Umwelt, 2015). This dataset consists of four simulations (4dp) (three for training and one for validation) and one larger simulation for testing scalability. Due to constraints of availability of subsurface measurements, see Figure˜1, the large-scale simulation is on a rectangular domain of only twice the length of the training data. Runtimes range from 38 to 91 hours (average 58 hours), and 134 hours for the scaling-test domain. Variation within the dataset for training stems from different heat pump placements in $i$ and from the location where $k$ is extract from the Munich region in 1.
Information on simulations and hydro-geological parameters are in Section˜A.1; hardware specifications in Section˜A.4.
@@
We evaluate model accuracy separately for each output field (directional velocities $v_{x},v_{y}$, and temperature $T$) using standard ML metrics of Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Huber loss; the structure-focused metric of Structural Similarity Index Measure (SSIM) (Wang et al., 2004), and application-driven metrics of Measure of Connectivity (MoC) and visual assessment. MoC expresses how physically reasonable the temperature fields are by measuring the percentage of predicted heat plume cells that are not connected via 4-connectivity [^1] This means axis-aligned neighbors only, excluding diagonal connections. to a heat pump based on a temperature threshold of $0.1$ °C, corresponding to measurement precision (UKB System Technology, 2024). All metrics other than SSIM are applied after re-normalization to the original data ranges to obtain physically meaningful results in [°C] or [meters/year].
3 Application of Established ML Approaches
To motivate a structure-specific architecture, we present the performance of six established ML approaches for predicting $T$ from $p,k,i$: a UNet, a domain decomposition-based UNet (DDUNet) (Verburg et al., 2025), a dilated UNet, UNet++ (Zhou et al., 2018), an FNO and a PINN (results in Tables˜1 and 2, parameters in Section˜A.2.3).
PINNs and FNOs
PINNs, while promising for low-data fluid dynamics tasks (Rao et al., 2020; Sun et al., 2020; Cai et al., 2021; Takamoto et al., 2022), struggle in preliminary tests with complex scenarios (Krishnapriyan et al., 2021), such as heterogeneous media and discontinuous source terms. Physics-informed CNNs (PI-CNNs) have been evaluated in simplified single-plume aquifer settings (Hirche, 2023; Pelzer & Piller, 2024) and consistently failed to outperform standard CNNs, even when the physics-based loss was used only as an auxiliary regularization term alongside a data loss. The largest errors occur near injection locations, likely due to discontinuous source term fields. The largest errors occur near injection locations, likely due to discontinuous source term fields. The largest errors occur near injection locations, likely due to discontinuous source term fields. In our 3dp problem, overfitting a PI-CNN to a single datapoint using only energy-based loss terms (cf. Section˜A.2.2) derived from the governing equations of the simulations fails to model $T$ but instead predicts a field that resembles $v_{y}$ (Figure˜2). This behavior is consistent with the model capturing only a preliminary stage of the physical process and indicates a representational limitation of the learning setup regarding transport modeling.
While FNOs are in principle capable of modeling both short- and long-range interactions, we found them impractical for the target problem due to infeasible memory requirements at realistic domain sizes and their sensitivity to non-periodic boundary conditions. Additionally, they performed poorly in the presence of multiple sources, consistent with Liu-Schiaffini et al. (2024). On our example, we had to reduce the domain size to $\times480$ cells to fit it on a graphics processing unit (GPU) (Section˜A.4) during training and still only managed to predict low frequency aspects, i.e., the background temperature, in our overfitting test to one datapoint.
UNet-based Methods
Convolutional architectures such as UNet primarily capture local spatial dependencies and can, in principle, generalize from a single large-domain sample when interactions are local. Modeling global effects, however, requires enlarged receptive fields or architectural adaptations, such as dilated convolutions. UNet++ fuses multiscale features through nested skip connections, while still relying on convolutional locality. DDUNet addresses large domains under GPU memory constraints by decomposing the domain into, here $2\times 2$, subdomains with learned coarse-scale feature exchange, enabling effective global context at reduced memory cost. The hyperparameters of all models are optimized according to Section˜A.2.3.
Table 1: Metrics of all experiments. Results for standard models and LGCNN (end-to-end and Steps 1 and 3) on synthetic and real permeability fields $k$, including ablation variants. Models are trained on 3dp, 101dp, or 4dp. Metrics are reported in physical units ([°C] for Step 3 and end-to-end, [m/y] for Step 1), except MoC ([%]) and SSIM (unitless). Velocity experiments are shaded due to differing scales; experiment details are referenced in column Sec. Only contains representative runs, statistical errors are reported in the appendix.
Architecture Training Data Applied to Step MAE RMSE Huber SSIM MoC Sec.
Standard UNet synth. 101dp test end-to-end 0.0473 0.1001 0.0048 0.8292 0.09 3
UNet++ synth. 101dp test end-to-end 0.0918 0.1435 0.0102 0.5331 4.06 3
DDUNet synth. 101dp test end-to-end 0.0548 0.1134 0.0063 0.8120 0.59 3
UNet synth. 3dp test end-to-end 0.1908 0.6190 0.0736 0.6671 10.85 3
Dilated UNet synth. 3dp test end-to-end 0.1200 0.2236 0.0237 0.5004 10.95 3
PI-CNN (overfit) synth. 1dp train end-to-end 0.7532 0.8593 0.3454 0.1793 - 3
Ours LGCNN synth. 3dp test end-to-end 0.0916 0.1738 0.0146 0.6841 0.09 5.1
UNet synth. 3dp test 3 0.0417 0.0762 0.0029 0.8540 0.07 5.1
UNet synth. 3dp test 1 (vx) 22.3178 31.1860 21.8237 0.9739 - 5.1
UNet synth. 3dp test 1 (vy) 32.7444 45.0703 32.2488 0.9733 - 5.1
Experiments UNet, in sequence synth. 3dp test 3 0.0919 0.1695 0.0139 0.6714 1.30 5.2
UNet, w/ zero-padding synth. 3dp test 3 0.0487 0.0795 0.0031 0.7784 0.57 5.2
UNet, w/o partitioning synth. 3dp test 3 0.0429 0.0793 0.0031 0.8349 0.29 5.2
UNet synth. 101dp test 3 0.0416 0.0816 0.0032 0.8016 0.47 5.2
DDUNet synth. 101dp test 3 0.0376 0.0764 0.0029 0.8843 0.26 5.2
UNet, vary inputs synth. 3dp test 3 cf. A.3 cf. A.3 cf. A.3 cf. A.3 cf. A.3 A.3
UNet, w/o partitioning synth. 3dp test 1 (vx) 23.8584 33.8297 23.3646 0.9765 - 5.2
UNet, w/o partitioning synth. 3dp test 1 (vy) 35.9249 49.9040 35.4289 0.9721 - 5.2
UNet synth. 101dp test 1 (vx) 12.7036 29.9379 12.2161 0.9437 - 5.2
UNet synth. 101dp test 1 (vy) 324.7751 328.4992 324.2751 0.9572 - 5.2
DDUNet synth. 101dp test 1 (vx) 13.6736 36.9852 13.1829 0.9468 - 5.2
DDUNet synth. 101dp test 1 (vy) 322.0676 325.6697 321.5676 0.9741 - 5.2
Scaling DDUNet (vanilla) synth. 101dp scaling end-to-end 0.0235 0.0723 0.0025 0.8874 1.18 3
UNet (vanilla) synth. 101dp scaling end-to-end 0.0202 0.0578 0.0016 0.8914 0.54 3
LGCNN synth. 3dp scaling end-to-end 0.0413 0.1189 0.0065 0.6511 0.06 5.1
UNet synth. 3dp scaling 3 0.0168 0.0373 0.0007 0.9405 0.08 5.1
UNet synth. 3dp scaling 1 (vx) 24.9261 34.7004 24.4314 0.9869 - 5.1
UNet synth. 3dp scaling 1 (vy) 26.2795 38.2599 25.7847 0.9932 - 5.1
Domain transfer LGCNN real 4dp val end-to-end 0.0841 0.1659 0.0137 0.6511 1.65 5.3
UNet real 4dp val 3 0.0175 0.0319 0.0005 0.9405 3.30 5.3
UNet real 4dp val 1 (vx) 13.3316 16.5833 12.8383 0.9869 - 5.3
UNet real 4dp val 1 (vy) 9.0224 10.5612 8.5327 0.9932 - 5.3
LGCNN real 4dp scaling end-to-end 0.0394 0.0665 0.0022 0.6755 3.50 5.3
UNet real 4dp scaling 3 0.0189 0.0287 0.0004 0.8942 1.77 5.3
UNet real 4dp scaling 1 (vx) 110.0078 116.4903 109.5079 0.9436 - 5.3
UNet real 4dp scaling 1 (vy) 17.6051 24.8298 17.1118 0.9905 - 5.3
UNet and dilated UNet are trained on 3dp; UNet, UNet++ and DDUNet are trained on 101dp directly (Section˜A.3). All models are evaluated on the 3dp test data for direct comparability to LGCNN, some on the scaling datapoint. Training on 3dp leads to short, disconnected plumes (Figure˜2) with high MoC. When trained on 101dp, UNet and DDUNet show robust predictions and reasonable scaling, with DDUNet occasionally hallucinating plumes but offering lower memory usage and shorter runtimes (see Table˜12). UNet++ fails due to memory constraints in the scaling test.
Overall, well-chosen established approaches can perform reasonably (best is UNet101dp: MAE = 0.05°C, MoC = 0.09) and some can scale to larger domains, but none performs reasonably well in the data-scarce setting we are interested in (best is UNet3dp: MAE = 0.19°C, MoC = 10.85). Their limits are clear: they are biased toward local patterns, require large datasets to capture long-range transport, and quickly run into compute and memory constraints at realistic scales. In low-data regimes they yield short, disconnected plumes or hallucinated structures, and scalability can break down (e.g., FNO and UNet++ fail under memory constraints).